In order to stop a 1500 kg car that is originally moving at 25 m/s, and if the car stops over a distance of 50 m, what average force must act on the car?
Answer #1

v^2 - u^2=2as
0-25*25=-2a*50
a=6.25m/s2

F=ma
F=1500*6.25
F=9375N
Answer #2

Avg accleration = U2/2s = 252/2*50 = 6.25 m/sec^2

So avg force = Ma = 1500*6.25 = 9375 N (ans)
Answer #3

Mo -A.xtNych.ct V u2 t2aA 62 S 2x SO 6.25 ms Vautat 6.2S aua 1300 2S-o 931S N